A family goes on vacation to a water park. During the drive there they averaged 48mph. On the return trip they averaged 80mph. If the distance each way was equal, and the whole trip driving was 6.4 hours, find the distance ONE WAY to the water park.
Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39632)

Travel Rates Rule, RT=D

                  rate         time        distance
GOING              48          d/48           d
BACK               80          d/80           d
Total                           6.4

Clear enough what to do


d=192

Answer by stanbon(75887)

A family goes on vacation to a water park. During the drive there they averaged 48mph. On the return trip they averaged 80mph. If the distance each way was equal, and the whole trip driving was 6.4 hours, find the distance ONE WAY to the water park.
---------
To the park DATA:
dist = x mph ; rate = 48mph ; time = d/r = x/48 hrs
------
From the park DATA:
dist = x mph ; rate = 80 mph ; time = x/80 hrs
--------
Equation:
time + time = 6.4 hrs
x/48 + x/80 = 6.4
------
80x + 48x = 6.4*80*48
------
128x = 24976
---
x = distance = 192 miles
----------------
Cheers,
Stan H.
------------